OpenStudy (anonymous):

art= 26% history=14% english=0.21 reading=0.12 math=? what percent chose math? write the percent as a decimal

OpenStudy (anonymous):

art=26% history=14% english=0.21=21/100=21% reading=0.12=12/100=12% \[math=100-(26+14+21+12)=youranswer\]

OpenStudy (anonymous):

i got 74

OpenStudy (anonymous):

\[math=100-(26+14+21+12)=100-73=27%\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

math=27%

OpenStudy (anonymous):

ok